# Car-Following Models: A Multidisciplinary Review

### v4

@@ II-A3 Psycho-physical Models
Action Points (AP)
Michaels and Cozen (1963) [94]
@@
2. Adaptability: Action Point model can be adapted to represent different driver behaviors, which is particularly effective in simulating different traffic components.
1. Complexity: Due to modeling multiple action points, it can be challenging for calibration. 2. Discontinuity: Discrete nature of action points can lead to abrupt changes.
Fuzzy Logic Models
Kikuchi and Chkkroborty (1992) [105]
@@
Todosiev [95] adopted a fundamental psycho-physical approach and investigated the thresholds where drivers change their behavior at an action point. Evans and Rothery [96] conducted experiments to quantify the thresholds between different phases. Wiedemann and Reiter [97] developed the well-known psycho-physical car-following model used in VISSIM. Another simulation program, PARAMICS, has incorporated the Fritzsche car-following model [98], an acceleration-based model from the psycho-physical car-following framework. As described in literature [99], once the absolute value of angular velocity exceeds its threshold, a driver notices that ego-vehicle speed is different from that of the leading vehicle and reacts with an ac/deceleration opposite in sign to that of the rate of change of visual angle. Winsum [100] integrated preferred time headway and Time-to-Collision (TTC) in a mathematical car-following model based on psychological evidence that human drivers regulate available time as a control mechanism. Wang et al. [101] devised ac/deceleration model based on the driver’s cognitive mechanism using the Just-Noticeable Distance (JND) concept. Lochrane et al. [102] developed a multidimensional framework for modeling thresholds to account for the different behavior in the work and nonwork zones. Durrani and Lee [103] calibrated the eight Calibration Constants (CCs) built into the Wiedemann model used by VISSIM to quantify the threshold values for drivers’ APs. Wagner et al. [104] analyzed a large dataset of car-following pairs to investigate the distribution of APs in terms of speed differences and gaps simulation with SUMO’s implementation of the Wiedemann model, finding that small distances indicate a slightly higher probability of finding an AP.
@@ II-B Adaptive Cruise Control Model
Unlike the previous models designed to simulate human driver behavior and reveal the disturbances of natural traffic, ACC models are developed to dampen the instability from the mechanistic view. ACC systems directly manage the basic operations of the vehicle, such as throttle control, braking, and steering. Low-level controllers are responsible for executing the commands to meet the higher-level objectives, like maintaining a set speed or following distance in ACC. Adaptive cruise control has been researched since the 1970s[111, 112, 113, 114, 115, 116] and was first on the market when Mitsubishi launched the ”Preview Distance Control” driver assistance system in 1995. ACC and CACC (Cooperative Adaptive Cruise Control) are heavily researched CAV applications. ACC models are designed for autonomous car following to reduce the feasible time headway among vehicles [117]. This paper [118] examines challenges of CACC systems in real-world deployment, including routing protocols, driver engagement, and adapting control designs to realistic traffic conditions. Some studies [119, 120, 121] suggest that the Constant Time Gap policy is more robust against error propagation through traffic than the Constant Distance Gap policy. Table IV summarizes representative adaptive cruise control models.
@@ II-B1 Linear Model
Linear controller assumes acceleration is proportional to the deviation from target spacing and relative speed. Helly [122] proposed the first linear model that prescribed the adaptation of the acceleration of each vehicle as a function of relative velocity, relative distance from the leading car, and other pertinent parameters. Milanés et al.[123] compared four control techniques: PI, intelligent PI, fuzzy controller, and adaptive-network-based fuzzy control, showing that the intelligent PI performs best. Dias et al.[125] used a linear longitudinal force model as a feedforward controller or Proportional-Integral-Derivative (PID) to compensate for the powertrain non-linearity. The LQ-based controller has been applied on ACC and CACC systems and tested under rural and urban roads in a CV environment, considering both communication delay and driver reaction time [124, 126]. A classical Constant-Time Headway Relative-Velocity (CTH-RV) model [127] with gains of speed error and spacing error is described as follows.
@@ II-B3 Model Predictive Control (MPC)
Linear Model
Helly (1959) [122]
@@
Nonlinear Model
Holve, et al. (1995) [129]
@@
Model Predictive Control
Bageshewar, et al. (2004) [135]
@@
III DATA-DRIVEN MODELS
@@
Machine Learning/ Deep Learning
Kehtarnavaz et al. (1998) [142]
@@
1. Data-driven: ML/DL models learn directly from data, allowing them to capture the complexity of real-world driving behavior without the need for explicit modeling assumptions.
2. Flexibility: A wide range of machine learning and deep learning techniques can be used, enabling the selection of an appropriate model for the specific problem and data.
3. Driver characteristics: Supervised learning models can learn to capture heterogeneous driver characteristics or driving styles by including driver- specific features in the input data.
1. Labeled data requirements: Supervised learning models require a large amount of labeled training data, which can be expensive and time-consuming to collect and annotate.
2. Cascading errors: In multistep prediction tasks, an incorrect estimation at an early stage can propagate through subsequent steps, leading to increasingly inaccurate decisions.
3. Interpretability and Explainability: Complex deep learning models, can be difficult to interpret and explain, which can be problematic in safety-critical applications.
Imitation Learning
@@
Unsupervised Learning
Koutsopoulos and Farah (2012) [167]
@@
Reinforcement Learning
Ioannou and Chien (1993) [171]
@@ III-A1 Machine Learning Model
With the availability of high-resolution vehicle trajectory data and the rampant development of machine learning (ML) algorithms and tools, several high-accuracy machine learning and deep learning algorithms have been developed. Kehtarnavaz et al.[142] developed a time-delay neural network (TDNN) autonomous vehicle following module to replace a conventional proportional and derivative/proportional and integral (PD/PI) controller and implemented it on actual vehicles. Panwai and Dia [143] developed a neural agent car-following model for mapping perceptions to actions. They compared the speed and position of individual vehicles from data to model output trajectory profiles. Khodayari et al. [144] developed an Artificial Neural Network (ANN) to simulate and predict the car-following behavior based on the driver-vehicle unit’s reaction delay, and input parameters include relative speed, relative distance, and follower’s speed. Zheng et al. [145] developed ANN to output the following vehicle’s speed by considering the driver–vehicle reaction delay in relative speed and acceleration, the gap, and the vehicle’s speed. Wei and Liu [146] proposed a Support Vector Machine approach to investigate the asymmetric characteristics of car-following behavior and its impact on traffic flow evolution. The Support Vector Machine model takes space headway, velocity, and relative speed as the inputs and the output is the follower’s speed. Another study [147] proposed a data-driven locally weighted regression approach for optimizing car-following model estimation by applying an existing technique and calibrated using instrumented vehicles in Naples, Italy. Cheng et al. [148] developed a hybrid method by integrating theoretical-based and machine-learning models to obtain better accuracy while maintaining the physical meaning. Kamjoo et al. [149] implemented Support Vector Machine and Long-Short-Term-Memory (LSTM) models to investigate the winter operation’s impact on car-following behavior, showing that the presence of snowplows leads to significantly different car-following behaviors.
@@ III-A2 Deep Learning Model
Driving behaviors can be treated as sequential data since the driver’s following action is conditioned on their previous action. The Deep Learning (DL) model, particularly Recurrent Neural Networks (RNN), is similar to humans in memory-based decision-making. Morton et al. [150] evaluated LSTM networks for simulating driver acceleration behavior on highways, which can effectively replicate real driver behavior in traffic models, matching or surpassing baseline methods in accuracy and realism. Zhou et al. [151] developed an RNN with inputs of the gap, relative speed, and follower’s speed to predict acceleration, speed, and gap that can be used for car-following control. Wang et al. (2017) [152] developed deep learning-based car-following with Gated Recurrent Unit (GRU) that embeds model outputs’ prediction and memory effects. Wu et al. (2019) [153] built a deep learning-based model to mimic human drivers’ memory, attention, and prediction (MAP) mechanisms. Wang et al. (2019) [154] analyzed the long-term memory effect in the car-following model using a deep learning model by taking various time-horizon historical information as inputs. Zhang et al. [155] built an LSTM model with a hybrid retraining constrained (HRC) training method to model car-following and lane-changing behaviors. Given that CF behaviors (acceleration, deceleration, and cruising) are made for continuous time steps rather than step-by-step with memory effects and delayed reaction time, Ma and Qu [156] developed a seq2seq LSTM model for predicting multistep car-following behaviors. Their model inputs include gap distance, relative speed difference, and the speed of the subject vehicle, and the model output is the multiple-step accelerations/decelerations. Mo et al. [157] designed Physic-Informed Deep Learning architectures encoded with IDM and OVM to predict accelerations in four traffic regimes: acceleration, deceleration, cruising, and emergency braking. Zhou et al. [158] applied a transfer learning-based LSTM car-following model for adaptive cruise control to address the ACC data scarcity problem by transferring useful features from human-driven data. Missing data in vehicle state information often leads to gaps in the data required for making accurate control decisions. A longitudinal CF model leverages Transformer-Generative Adversarial Networks (TransGAN) [159] for this purpose, which is particularly significant in scenarios where data acquisition devices malfunction.
@@ III-A3 Imitation Learning Models
Imitation learning (IL) a subset of machine learning focused on learning from demonstration, which has several advantages: 1. It can meet the preferences of driving styles; 2. Other drivers would easily understand the automated car-following maneuver. Behavior Cloning (BC) models, which are the simplest forms of Imitation Learning, try to learn the maneuver of the human driver given the expert’s demonstrations that are divided into state-action pairs. The BC model can discover complex car-following strategies for nonlinear and high-dimensional systems. Inverse reinforcement learning (IRL) can be seen as a form of imitation learning where the model is not just learning to mimic the actions of the expert, but is actually trying to understand the motivations (i.e., the reward function) behind those actions. This can lead to more robust and versatile models, as they are not just copying behavior, but are learning the principles that guide that behavior. Kuefler et al. [163] proposed Generative Adversarial Imitation Learning (GAIL) framework to imitate human car-following behavior with three non-linearity in mapping from states to actions, high-dimensional state representation, and stochasticity. Instead of learning a policy from a reward function, Inverse Reinforcement Learning (IRL) is trying to learn a reward function from the demonstration of a task [160]. With naturalistic driving behavior data, the IRL would approximate that human’s reward function for the driving task. Shimosaka et al. [161] proposed an IRL framework to model risk anticipation and defensive driving, which has a better long-term prediction of driver maneuvers. Shimosaka et al. [162] developed multiple reward functions with clustering of environments to better handle environmental diversity. The framework then simultaneously learns the parameters specific to each cluster. This method significantly improves the model’s adaptability and accuracy in diverse settings. Gao et al. [164] proposed a car-following algorithm with a reward function based on IRL under different conditions. The experimental verification is conducted on a dynamic driving simulation test bench. Zhou [165] developed a Maximum Entropy Deep IRL approach using a neural network to estimate the rewards. Model input features include time headway, relative speed, and maximum speed, and model output is speed control. Zhao et al. [166] developed a Personalized-ACC framework using IRL, which classified the driver and weather conditions for the pre-trained off-line IRL model and compared it with the IDM model regarding safety and comfort measured by takeover percentage.
@@ III-C Reinforcement Learning
Ioannou and Chien [171] applied RL to develop an autonomous intelligent adaptive cruise control system. Desjardins and Chaib-draa [172] introduced a policy gradient algorithm estimation and used BPNN for longitudinal vehicle control. Their RL-based car-following control was implemented in the simulator to realize the efficient behavior of CACC. Zhu et al. [173] proposed a framework for developing human-like car-following based on deep reinforcement learning using speed deviations as a reward function. Zhang et al. [174] developed a deterministic promotion RL algorithm (DPRL) based on an actor-critic frame and the policy gradient method for longitudinal velocity control, implemented in CarSim simulated environment and real-world experiment. Li et al. [175] proposed a deep deterministic policy gradient (DDPG)-based driving strategy taking information from in-vehicle sensors for system performance optimization. Yen et al. [176] developed a proactive car-following model using deep-RL considering road efficiency, safety, and comfort, which shows better performances in terms of time heady, time to collision (TTC), and jerk (rate of change in acceleration). Yavas et al. [177] developed a Model-Based Reinforcement Learning (MBRL) for ACC control with a hybrid policy that combines the Intelligent Driver Model following policy with the Deep Reinforcement Learning policy. Wang et al. [178] devised an RL model for longitudinal velocity control in a three-vehicle mode based on the safety level, consisting of the collision avoidance priority of the leading and following vehicles and the expected acceleration/deceleration decision. Most CF models take relative speed, acceleration, and gap as inputs, while some studies [179, 180] use video object detection to learn the driving strategy. Such models demonstrate promising results, which can efficiently be deployed on embedded devices for advanced driver assistance systems.
IV DISCUSSION
IV-A Applications

### v5

@@ II-A3 Psycho-physical Models
Action Points (AP) Michaels and Cozen (1963) [100]
@@
2. Adaptability: Action Point model can be adapted to represent different driver behaviors, which is particularly effective in simulating a variety of traffic components.
1. Complexity: Due to modeling multiple action points, it can be challenging for calibration.
2. Discontinuity: Discrete nature of action points can lead to abrupt changes.
Fuzzy Logic Models Kikuchi and Chkkroborty (1992) [111]
@@
Todosiev [101] adopted a fundamental psycho-physical approach and investigated the thresholds where drivers change their behavior at an action point. Evans and Rothery [102] conducted experiments to quantify the thresholds between different phases. Wiedemann and Reiter [103] developed the well-known psycho-physical car-following model used in VISSIM. Another simulation program, PARAMICS, has incorporated the Fritzsche car-following model [104], an acceleration-based model from the psycho-physical car-following framework. As described in literature [105], once the absolute value of angular velocity exceeds its threshold, a driver notices that ego-vehicle speed is different from that of the leading vehicle and reacts with an ac/deceleration opposite in sign to that of the rate of change of visual angle. Winsum [106] integrated preferred time headway and Time-to-Collision (TTC) in a mathematical car-following model based on psychological evidence that human drivers regulate available time as a control mechanism. Wang et al. [107] devised ac/deceleration model based on the driver’s cognitive mechanism using the Just-Noticeable Distance (JND) concept. Lochrane et al. [108] developed a multidimensional framework for modeling thresholds to account for the different behavior in the work and nonwork zones. Durrani and Lee [109] calibrated the eight Calibration Constants (CCs) built into the Wiedemann model used by VISSIM to quantify the threshold values for drivers’ APs. Wagner et al. [110] analyzed a large dataset of car-following pairs to investigate the distribution of APs in terms of speed differences and gaps with SUMO’s implementation of the Wiedemann model, finding that small distances indicate a slightly higher probability of finding an AP.
@@ II-B Adaptive Cruise Control Model
Unlike the previous models designed to simulate human driver behavior and reveal the disturbances of natural traffic, ACC models are developed to dampen the instability from the mechanistic view. ACC systems directly manage the basic operations of the vehicle, such as throttle control, braking, and steering. Low-level controllers are responsible for executing the commands to meet the higher-level objectives, like maintaining a set speed or following distance in ACC. Adaptive cruise control has been researched since the 1970s[117, 118, 119, 120, 121, 122] and was first on the market when Mitsubishi launched the ”Preview Distance Control” driver assistance system in 1995. ACC and CACC (Cooperative Adaptive Cruise Control) are heavily researched CAV applications. ACC models are designed for autonomous car following to reduce the feasible time headway among vehicles [123]. This paper [124] examines challenges of CACC systems in real-world deployment, including routing protocols, driver engagement, and adapting control designs to realistic traffic conditions. Some studies [125, 126, 127] suggest that the Constant Time Gap policy is more robust against error propagation through traffic than the Constant Distance Gap policy. By taking into account underlying the vehicle dynamics, e.g., force, torque, and power, Rakha et al. [128, 129] combined vehicle powertrain model with driver behavior to realize more realistic modeling of intelligent vehicle applications. Fadhloun-Rakha (FR) model [130, 131] was proposed by capturing heterogeneous driving behavior and vehicle engine dynamics, considering two constraints regarding car following interactions (collision avoidance and steady-state traffic stream). Table IV summarizes representative adaptive cruise control models.
@@ II-B1 Linear Model
Linear controller assumes acceleration is proportional to the deviation from target spacing and relative speed. Helly [132] proposed the first linear model that prescribes the adaptation of the acceleration of each vehicle as a function of relative velocity, relative distance from the leading car, and other pertinent parameters. Milanés et al.[133] compared four control techniques: PI, intelligent PI, fuzzy controller, and adaptive-network-based fuzzy control, showing that the intelligent PI performs best. Dias et al.[135] used a linear longitudinal force model as a feedforward controller or Proportional-Integral-Derivative (PID) to compensate for the powertrain non-linearity. The LQ-based controller has been applied on ACC and CACC systems and tested under rural and urban roads in a CV environment, considering both communication delay and driver reaction time [134, 136]. A classical Constant-Time Headway Relative-Velocity (CTH-RV) model [137] with gains of speed error and spacing error is described as follows.
@@ II-B3 Model Predictive Control (MPC)
Linear Model Helly (1959) [132]
@@
Nonlinear Model Holve, et al. (1995) [139]
@@
Model Predictive Control Bageshewar, et al. (2004) [145]
@@
III Learning-based Models
@@
Machine Learning Kehtarnavaz et al. (1998) [152]
@@
Dabiri and Abbas (2018) [158]
Yang et al. (2018) [159]
Soldevila et al. (2021) [160]
Kamjoo et al. (2022) [161]
1. Interpretability: Traditional ML models (e.g. Linear Regression, Decision Tree and SVM) are interpretable, making it easier to understand how decisions are being made, which is crucial for safety-critical applications.
2. Faster Calibration and Inference: Due to their simpler nature, traditional ML models can be trained faster and often have lower latency during inference.
1. Feature Engineering: ML models often rely heavily on manual feature engineering.
2. Scalability: As the complexity of the scenario increases, the performance of traditional ML models might not scale as well as other models.
Deep Learning Morton et al. (2016) [162]
@@
Lu et al. (2023) [171]
Xu et al. (2024) [172]
Chen et al. (2024) [173]
1. Data-driven: DL models learn directly from data, allowing them to capture the complexity of real-world driving behavior without the need for explicit modeling assumptions.
2. Flexibility: A wide range of DL techniques can be used, enabling the selection of an appropriate model for the specific problem and data.
3. Incorporation of driver characteristics: DL models can learn to capture driver characteristics or latent driving styles, enabling the modeling of heterogeneous human driving behaviors.
1. Labeled data requirements: Supervised learning models require a large amount of labeled training data.
2. Cascading errors: In sequential prediction tasks, an incorrect estimation at an early stage can propagate through subsequent steps, leading to increasingly inaccurate decisions.
Imitation Learning Shimosaka et al. (2014) [175]
@@
Unsupervised Learning Koutsopoulos and Farah (2012) [181]
@@
Reinforcement Learning Ioannou and Chien (1993) [185]
@@ III-A1 Machine Learning Model
With the availability of high-resolution vehicle trajectory data and the rampant development of machine learning (ML) algorithms and tools, several high-accuracy machine learning and deep learning algorithms have been developed. Kehtarnavaz et al. [152] developed a time-delay neural network (TDNN) autonomous vehicle following module to replace a conventional proportional and derivative/proportional and integral (PD/PI) controller and implemented it on actual vehicles. Panwai and Dia [153] developed a neural agent car-following model for mapping perceptions to actions. They compared the speed and position of individual vehicles from data to model output trajectory profiles. Khodayari et al. [154] developed an Artificial Neural Network (ANN) to simulate and predict the car-following behavior based on the driver-vehicle unit’s reaction delay, and input parameters include relative speed, relative distance, and follower’s speed. Zheng et al. [155] developed ANN to output the following vehicle’s speed by considering the driver–vehicle reaction delay in relative speed and acceleration, the gap, and the vehicle’s speed. Wei and Liu [156] proposed a Support Vector Machine approach to investigate the asymmetric characteristics of car-following behavior and its impact on traffic flow evolution. The Support Vector Machine model takes space headway, velocity, and relative speed as the inputs and the output is the follower’s speed. Another study [157] proposed a data-driven locally weighted regression approach for optimizing car-following model estimation by applying an existing technique and calibrated using instrumented vehicles in Naples, Italy. The gradient boosting of regression tree (GBRT) algorithm [158] was applied to NGSIM data to develop a car-following model that outperforms traditional GHR models in capturing the motion characteristics of successive vehicles, using cross-validation and sensitivity analysis for parameter tuning. Yang et al. [159] developed a hybrid method by integrating theoretical-based and machine-learning models to obtain better accuracy while maintaining the physical meaning. Kamjoo et al. A Gaussian Process Regression Approach [160] to model individual longitudinal driving behaviors by integrating parametric and non-parametric methods, enhancing predictions with large data sets and adapting to new variables without changing the model structure. [161] implemented Support Vector Machine and Long-Short-Term-Memory (LSTM) models to investigate the winter operation’s impact on car-following behavior, showing that the presence of snowplows leads to significantly different car-following behaviors.
@@ III-A2 Deep Learning Model
Driving behaviors can be treated as sequential data since the driver’s following action is conditioned on their previous action. Deep Learning (DL) models, particularly Recurrent Neural Networks (RNN), are similar to humans in memory-based decision-making. Morton et al. [162] evaluated LSTM networks for simulating driver acceleration behavior on highways, which can effectively replicate real driver behavior in traffic models, matching or surpassing baseline methods in accuracy and realism. Zhou et al. [163] developed an RNN with inputs of the gap, relative speed, and follower’s speed to predict acceleration, speed, and gap that can be used for car-following control. Wang et al. (2017) [164] developed deep learning-based car-following with Gated Recurrent Unit (GRU) that embeds model outputs’ prediction and memory effects. Wang et al. (2019) [165] applied Gated Recurrent Unit (GRU) with different timescale historical information as inputs to test memory effects on car following behavior. Wu et al. (2019) [166] built a deep learning-based model to mimic human drivers’ memory, attention, and prediction (MAP) mechanisms. Zhang et al. [167] built an LSTM model with a hybrid retraining constrained (HRC) training method to model car-following and lane-changing behaviors. Given that CF behaviors (acceleration, deceleration, and cruising) are made for continuous time steps rather than step-by-step with memory effects and delayed reaction time, Ma and Qu [168] developed a seq2seq LSTM model for predicting multistep car-following behaviors. Their model inputs include gap distance, relative speed difference, and the speed of the subject vehicle, and the model output is the multiple-step accelerations/decelerations. Zhou et al. [169] applied a transfer learning-based LSTM car-following model for adaptive cruise control to address the ACC data scarcity problem by transferring useful features from human-driven data. Missing data in vehicle state information often leads to gaps in the data required for making accurate control decisions. A longitudinal CF model leverages Transformer-Generative Adversarial Networks (TransGAN) [170] for this purpose, which is particularly significant in scenarios where data acquisition devices malfunction. An improved sequence-to-sequence model for connected automated vehicles (CAVs) [171] enhances the accuracy of simulating speeds and positions by incorporating kinematics information from multiple preceding vehicles and using a novel deep learning framework. Another sequence-to-sequence model [172] incorporates an attention mechanism to learn the probability distribution of driver reaction delay. Aggressiveness Informed Car-Following (AICF) [173] was proposed approach, which integrates driving style as a dynamic input in data-driven car-following models to address the limitations of using predefined patterns and categories.
@@ III-A3 Imitation Learning Models
Imitation learning (IL) a subset of machine learning focused on learning from demonstration, which has several advantages: 1. It can meet the preferences of driving styles; 2. Other drivers would easily understand the automated car-following maneuver. Behavior Cloning (BC) models, which are the simplest forms of Imitation Learning, try to learn the maneuver of the human driver given the expert’s demonstrations that are divided into state-action pairs. The BC model can discover complex car-following strategies for nonlinear and high-dimensional systems. Inverse reinforcement learning (IRL) can be seen as a form of imitation learning where the model is not just learning to mimic the actions of the expert, but is actually trying to understand the motivations (i.e., the reward function) behind those actions. This can lead to more robust and versatile models, as they are not just copying behavior, but are learning the principles that guide that behavior. With naturalistic driving behavior data, the IRL would approximate that human’s reward function for the driving task. Instead of learning a policy from a reward function, Inverse Reinforcement Learning (IRL) is trying to learn a reward function from the demonstration of a task [174]. Shimosaka et al. [175] proposed an IRL framework to model risk anticipation and defensive driving, which has a better long-term predict [176] developed multiple reward functions with clustering of environments to better handle environmental diversity. The framework then simultaneously learns the parameters specific to each cluster. This method significantly improves the model’s adaptability and accuracy in diverse settings. Kuefler et al. [177] proposed Generative Adversarial Imitation Learning (GAIL) framework to imitate human car-following behavior with three non-linearity in mapping from states to actions, high-dimensional state representation, and stochasticity. Gao et al. [178] proposed a car-following algorithm with a reward function based on IRL under different conditions. The experimental verification is conducted on a dynamic driving simulation test bench. Zhou [179] developed a Maximum Entropy Deep IRL approach using a neural network to estimate the rewards. Model input features include time headway, relative speed, and maximum speed, and model output is speed control. Zhao et al. [180] developed a Personalized-ACC framework using IRL, which classified the driver and weather conditions for the pre-trained off-line IRL model and compared it with the IDM model regarding safety and comfort measured by takeover percentage.
@@ III-C Reinforcement Learning
Ioannou and Chien [185] applied RL to develop an autonomous intelligent adaptive cruise control system. Desjardins and Chaib-draa [186] introduced a policy gradient algorithm estimation and used BPNN for longitudinal vehicle control. Their RL-based car-following control was implemented in the simulator to realize the efficient behavior of CACC. Zhu et al. [187] proposed a framework for developing human-like car-following based on deep reinforcement learning using speed deviations as a reward function. Zhang et al. [188] developed a deterministic promotion RL algorithm (DPRL) based on an actor-critic frame and the policy gradient method for longitudinal velocity control, implemented in CarSim simulated environment and real-world experiment. Li et al. [189] proposed a deep deterministic policy gradient (DDPG)-based driving strategy taking information from in-vehicle sensors for system performance optimization. Yen et al. [190] developed a proactive car-following model using deep-RL considering road efficiency, safety, and comfort, which shows better performances in terms of time heady, time to collision (TTC), and jerk (rate of change in acceleration). Yavas et al. [192] developed a Model-Based Reinforcement Learning (MBRL) for ACC control with a hybrid policy that combines the Intelligent Driver Model following policy with the Deep Reinforcement Learning policy. Wang et al. [193] devised an RL model for longitudinal velocity control in a three-vehicle mode based on the safety level, consisting of the collision avoidance priority of the leading and following vehicles and the expected acceleration/deceleration decision. Most CF models take relative speed, acceleration, and gap as inputs, while some studies [191, 194] use video object detection to learn the driving strategy. Such models demonstrate promising results, which can efficiently be deployed on embedded devices for advanced driver assistance systems. This study [195] develops a personalized car-following model using a memory-based deep reinforcement learning approach, combining Twin Delayed Deep Deterministic Policy Gradients (TD3) with Long Short-Term Memory (LSTM) networks to optimize car-following behavior based on diverse driving styles for improved safety, efficiency, and comfort, demonstrating superior performance and convergence compared to other models.
IV Knowledge-driven Models
Knowledge-driven methods leverage extensive pre-trained Large GenAI models and foundational knowledge to generate new content and solutions, simulating commonsense-based understanding and reasoning that reflect deep concepts. Large GenAI models can address the heterogeneity in traffic components, varying road environments, and the stochastic nature of driving with an agent-based approach. CF behavior is considered as a subtask of Autonomous Driving (AD), which suggests that large GenAI models developed in AD can also be applicable in CF applications. The kinematic CF models (e.g., IDM models) are useful as baseline for validating new knowledge-based driving agents. In CF scenarios, these knowledge-driven methods reason about vehicle behaviors, enhancing decision-making in complex driving scenarios. Additionally, they provide context-aware systems that understand and react to real-time driving conditions, leading to more realistic simulations, optimal control, and enhanced human-machine interaction.
IV-A Large Language Models (LLMs)
LLMs can generate detailed driving instructions and responses based on signal inputs and human-like understanding. These models can create complex driving policies by predicting and generating the next action based on previous driving scenarios and contextual data, which enhances the decision-making processes in autonomous CF, making vehicles more adaptive and responsive to dynamic driving environments. Examples of LLM-based autonomous driving models include DriveLLM [196], DiLu [198], Driving with LLMs [197], and Drive-as-you-speak [199].
IV-B Vision-Language Models (VLMs)
VLM models integrate visual inputs, such as images and videos, with textual descriptions to enhance understanding and generate new insights or content. These models excel in scenarios that require multi-modal perception and comprehension, enabling accurate interpretations of complex driving situations. Representative models include vision-language-action driving model (VLAM) (LINGO-1 [200] & LINGO-2 [201]) and DriveVLM [202].
IV-C Multimodal Large Language Models (MLLMs)
MLLMs extend the capabilities of traditional LLMs and VLMs by integrating multiple types of data inputs. These models can process and understand a combination of various sensor input to provide a holistic view of the driving environment. A recent survey paper [203] on Multimodal Large Language Models for AD sheds light on the latest advancements and research directions in the field. Representative models include DriveMLM [204] and LMDrive [205].
V Discussion
In this review paper, the focus on master models (Figure 3) offers a structured approach by providing a clear framework for categorization and understanding. The Figure 4 illustrates the efficacy of these master models in delineating the progression and branching out of various derivative algorithms within the field. By doing so, it highlights the master models’ role as pivotal reference points from which the landscape of car-following behavior can be surveyed. Automatic literature tools [206] adeptly navigate the extensive body of literature to surface the most relevant studies, demonstrating the significant advantage of leveraging structured categorizations to manage and interpret the wealth of research in this evolving area of traffic dynamics and control systems. This paper’s focused review through the lens of multi-disciplines thus not only consolidates existing knowledge but also guides future investigations.
Figure 3: The Historical Timeline of Reviewed Master Models and Key Variants.
Figure 4: Navigating the Literature with Reviewed Master Models Using OVM and IDM as Examples
V-A Theory-Based vs. Learning-Based vs. Knowledge-Driven
Decades of theory-based driving behavior research are now merging with learning-based and knowledge-driven opportunities. Although the data-driven models will undoubtedly become critically important, the traditional driver behavior expertise remains relevant. It is auspicious to integrate rule-based models with the fast-evolving AI models, which can reason, learn, and fill in missing information to achieve better interpretability, generalizability, and reliability.
TABLE VI: Comparative Overview of Theory-Based Models vs Learning-Based Models vs Knowledge-Driven Models
Theory-Based Models Learning-Based Models Knowledge-Driven Models Strengths Strengths Strengths Solid Theoretical Foundation: Built on well-established principles and theories, offering predictable behavior within their domain of validity. Interpretability: Due to their reliance on explicit formulas and parameters, these models provide clear insights into the cause-and-effect relationships within the modeled phenomena. Generalizability: Capable of generalizing across different scenarios, ensuring that the model adheres to expected physical laws and behavioral norms, which helps in maintaining performance even under unusual conditions not covered by the data. Accuracy and Prediction: Learning-Driven models are adept at capturing non-linear patterns from large datasets. This can lead to highly accurate predictions based on real-world data. Adaptability to Complex Environments: Ability to achieve human-level or superior performances under diverse driving conditions. Handling of Heterogeneous Data: Efficient processing and learning from varied data inputs, adapting to different traffic and infrastructure conditions. Physical-Social Intelligence: They understand both physical laws and drivers’ expectations and emotional needs, enabling them to identify and comprehend various patterns, rules, and relationships in driving behavior. Handling Edge Cases: Knowledge-Driven Large GenAI models address edge cases and unusual scenarios by leveraging commonsense reasoning and proactive decision-making. Context-Aware Systems: They enable context-aware systems that understand and react to real-time driving conditions, enhancing decision-making, human-machine interaction, and the overall safety and efficiency of intelligent vehicles. Limitations Limitations Limitations Complexity of Traffic Dynamics: Struggles to account for all heterogeneous traffic compositions, infrastructure conditions, and varying data inputs. Parametric Restrictions: Limited by the need to pre-define parameters, which may not capture the full spectrum of driver behaviors or environmental conditions. Overfitting and Scalability: Due to environmental diversity and sensor variability, learning-based techniques perform well on the data on which they were trained but not always adaptable on new, unseen datasets or in more complex environments. Learning Undesirable Behaviors: Potential to learn from bad behaviors due to drivers’ unpredictability or poor data quality. Alignment Challenges: Knowledge-Driven pre-trained foundational models are end-to-end systems. Misalignment occurs when objectives and behavior do not align with the intended use case, leading to undesired consequences or vulnerabilities. Safety and Ethical Concerns: Addressing potential biases, prioritize safety, and ensuring transparency in decision-making, are ongoing challenges that need careful consideration.
Several hybrid CF models were developed by integrating rule-based and AI-based models, which leverages the strengths of different approaches, enhancing accuracy, robustness, and safety in systems like autonomous vehicle control, where both predictable behavior and adaptability to real-world complexity are crucial. Li et al. [207] utilized the adaptive Kalman filter algorithm to integrate the LSTM data-driven model with the IDM theoretical-driven model to build the car-following model. Mo et al. [208] designed Physic-Informed Deep Learning architectures encoded with IDM and OVM to predict accelerations in four traffic regimes: acceleration, deceleration, cruising, and emergency braking. This paper [209] focused on analyzing the stochastics and heterogeneity in traffic dynamics using a data-driven car-following model. It examined the shape of the distribution of a stochastic term in the car-following model, predicting accelerations accurately and identifying relationships between predicted distributions’ modality, standard deviation, and traffic states. This research [210] introduces the IDM-Bi-LSTM combination car-following model, which integrates the theory-driven Intelligent Driver Model (IDM) with a data-driven model optimized by Particle Swarm Optimization (PSO) to enhance trajectory prediction accuracy and maintain controllability and security. Numerical simulations demonstrate that this hybrid model significantly reduces prediction errors and outperforms both individual IDM and PSO-Bi-LSTM models in simulating heterogeneous and asymmetric driving behaviors.
The comparison emphasizes (Table VI) the contrast between Rule-based and AI-based approaches in modeling car-following behavior. While theory-based models offer a clear, predictable framework grounded in established principles, their limitations in handling the complexity and variability of real-world driving scenarios highlight the potential benefits of incorporating learning and commonsense knowledge. Conversely, AI-driven models’ ability to adapt to and learn from diverse environments is tempered by challenges related to interpretability, the quality of learned behavior, and the handling of rare but critical situations. Balancing these strengths and limitations often results in more reliable, accurate, and practical systems for real-world applications.
V-B Applications